# Second- and third-order properties of multidimensional Langevin equations

### 25 Dec 2023 v5

@@ 6.2. One-dimensional case
Apparently, there are 6 parameters ($A$, $D$, $a$, $b$, $\mu$, $\nu$); however, there are actually only 5 (quadratic drift has only 3 coefficients). Thus, if $\mu$ and $\nu>0$ are the actual mean and variance of $x$, then given these together with $D>0$ and $b$, the possible values of the pair $(A,a)$ must lie along a one-dimensional manifold. Thus in Eq. (81), $\mathbf{C}$ should be interpreted as a “nominal” value not necessarily corresponding to the actual covariance matrix.
@@
$\qquad{}+a\mathcal{O}(a,b){He}_{3}(x)+a\mathcal{O}((a,b)^{2}){He}_{4}(x)+\cdots,$
@@
$\qquad{}+(2a+a^{2}\mathcal{O}(a,b)){He}_{3}(x)+a\mathcal{O}(a,b){He}_{4}(x)+\cdots,$
@@
We now present some numerical simulations to confirm the above results. We simulate a one-dimensional stochastic process where $\langle\dot{x}\mid x\rangle$ follows Eq. (123) with $x\to x+a$ and inhomogeneous diffusion $(1+bx/2)^{2}$. We use an Euler–Maruyama discretization with time-step $\Delta t=0.05$ and trajectory length $10^{6}$ time-units. Although corrections to the third-order covariance functions are only cubic order in $a,b$, they turn out to be quite large, as shown in Fig. 5 (the exception being when $a=0$). The second-order covariance function is also affected, although less so (Fig. 6), but the deviation is still larger than what one might expect for a quadratic correction. However, the deviations are limited to approximate multiplication by a constant, and thus nonlinearity cannot be effectively discerned from the second-order covariance functions, as expected. For larger values of $|a|$, the covariance functions are nowhere near the theoretically predicted values.
@@ 6.5. Integrated variables
We can also change coordinates to ${x^{\prime}}^{i}:=f_{1}^{i}(\mathbf{x})$, where
@@ 7.3. Inhomogeneous diffusion
(227) $\left\langle{\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}\mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\end{pmatrix}\mid\mathbf{x}}\right\rangle=\begin{pmatrix}\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&&\\ \mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&&\\ \widehat{\mathbf{A}}^{(2)+}_{+}&\widehat{\mathbf{A}}^{(2)+}_{-}&\widehat{\mathbf{A}}^{(2)+}_{(2)+}&\widehat{\mathbf{A}}^{(2)+}_{(2)-}\\ \widehat{\mathbf{A}}^{(2)-}_{+}&\widehat{\mathbf{A}}^{(2)-}_{-}&\widehat{\mathbf{A}}^{(2)-}_{(2)+}&\widehat{\mathbf{A}}^{(2)-}_{(2)-}\end{pmatrix}\begin{pmatrix}\mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\end{pmatrix}$
To have time-reversibility of the covariance functions, we first need that:
@@
(231) ${{}\widehat{\mathbf{A}}^{n}}^{(2)\pm}_{\mp}\langle\mathbf{x}^{\mp}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{(2)\pm}_{(2)\mp}\langle\mathbf{x}^{(2)\mp}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle$ $=-\langle\mathbf{x}^{(2)\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle{{\mathbf{A}^{n}}^{\mp}_{\pm}}^{\mathsf{T}}$
@@
(233) $\left\langle{\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\\ \mathbf{x}^{+}\Delta y\\ \mathbf{x}^{-}\Delta y\end{pmatrix}\mid\mathbf{x}}\right\rangle=\begin{pmatrix}0&&&&&&\\ &\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&&&&\\ &\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&&&&\\ &\widehat{\mathbf{A}}^{(2)+}_{+}&\widehat{\mathbf{A}}^{(2)+}_{-}&\widehat{\mathbf{A}}^{(2)+}_{(2)+}&\widehat{\mathbf{A}}^{(2)+}_{(2)-}&&\\ &\widehat{\mathbf{A}}^{(2)-}_{+}&\widehat{\mathbf{A}}^{(2)-}_{-}&\widehat{\mathbf{A}}^{(2)-}_{(2)+}&\widehat{\mathbf{A}}^{(2)-}_{(2)-}&&\\ \widehat{\mathbf{A}}^{y(1)+}_{(0)}&\widehat{\mathbf{A}}^{y(1)+}_{+}&\widehat{\mathbf{A}}^{y(1)+}_{-}&\widehat{\mathbf{A}}^{y(1)+}_{(2)+}&\widehat{\mathbf{A}}^{y(1)+}_{(2)-}&\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}\\ \widehat{\mathbf{A}}^{y(1)-}_{(0)}&\widehat{\mathbf{A}}^{y(1)-}_{+}&\widehat{\mathbf{A}}^{y(1)-}_{-}&\widehat{\mathbf{A}}^{y(1)-}_{(2)+}&\widehat{\mathbf{A}}^{y(1)-}_{(2)-}&\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}\end{pmatrix}\begin{pmatrix}1\\ \mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\\ \mathbf{x}^{+}\Delta y\\ \mathbf{x}^{-}\Delta y\end{pmatrix}$
@@
(236) ${{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{\mp}\langle\mathbf{x}^{\mp}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{y(1)\pm}_{(2)\mp}\langle\mathbf{x}^{(2)\mp}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle$ $=\langle\mathbf{x}^{\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)\mp}_{\pm}}^{\mathsf{T}}+\langle\mathbf{x}^{\pm}{\mathbf{x}^{(2)\pm}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)\mp}_{(2)\pm}}^{\mathsf{T}}$
Again, the base case $n=1$ of Eq. (236) is a consequence of the aforementioned conditions, while Eq. (235)–(236) can be shown by mathematical induction on $n$.
@@
(237) $\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\\ \langle\mathbf{x}^{(2)+}\Delta y\rangle\\ \langle\mathbf{x}^{(2)-}\Delta y\rangle\end{pmatrix}=\begin{pmatrix}0&&&&\\ {\widehat{\mathbf{A}}}^{y(1)+}_{(0)}&\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&&\\ {\widehat{\mathbf{A}}}^{y(1)-}_{(0)}&\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&&\\ {\widehat{\mathbf{A}}}^{y(2)+}_{(0)}&\widehat{\mathbf{A}}^{(2)+}_{+}&\widehat{\mathbf{A}}^{(2)+}_{-}&\widehat{\mathbf{A}}^{(2)+}_{(2)+}&\widehat{\mathbf{A}}^{(2)+}_{(2)-}\\ {\widehat{\mathbf{A}}}^{y(2)-}_{(0)}&\widehat{\mathbf{A}}^{(2)-}_{+}&\widehat{\mathbf{A}}^{(2)-}_{-}&\widehat{\mathbf{A}}^{(2)-}_{(2)+}&\widehat{\mathbf{A}}^{(2)-}_{(2)-}\end{pmatrix}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\\ \langle\mathbf{x}^{(2)+}\Delta y\rangle\\ \langle\mathbf{x}^{(2)-}\Delta y\rangle\end{pmatrix}$
@@
Once again, the case $n=1$ for Eq. (239) is a consequence of the aforementioned conditions, while Eq. (238)–(239) can be shown assuming the case $n=1$.
@@ 7.4. Nonlinear drift
(240) $\left\langle{\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\\ \mathbf{x}^{(3)+}\\ \mathbf{x}^{(3)-}\\ \vdots\end{pmatrix}\mid\mathbf{x}}\right\rangle=\begin{pmatrix}0&&&&&&&\\ &\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&{\widehat{\mathbf{A}}}^{+}_{(2)+}&{\widehat{\mathbf{A}}}^{+}_{(2)-}&&&\\ &\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&{\widehat{\mathbf{A}}}^{-}_{(2)+}&{\widehat{\mathbf{A}}}^{-}_{(2)-}&&&\\ &{\widehat{\mathbf{A}}}^{(2)+}_{+}&{\widehat{\mathbf{A}}}^{(2)+}_{-}&{\widehat{\mathbf{A}}}^{(2)+}_{(2)+}&{\widehat{\mathbf{A}}}^{(2)+}_{(2)-}&{\widehat{\mathbf{A}}}^{(2)+}_{(3)+}&{\widehat{\mathbf{A}}}^{(2)+}_{(3)-}&\\ &{\widehat{\mathbf{A}}}^{(2)-}_{+}&{\widehat{\mathbf{A}}}^{(2)-}_{-}&{\widehat{\mathbf{A}}}^{(2)-}_{(2)+}&{\widehat{\mathbf{A}}}^{(2)-}_{(2)-}&{\widehat{\mathbf{A}}}^{(2)-}_{(3)+}&{\widehat{\mathbf{A}}}^{(2)-}_{(3)-}&\\ {\widehat{\mathbf{A}}}^{(3)+}_{(0)}&{\widehat{\mathbf{A}}}^{(3)+}_{+}&{\widehat{\mathbf{A}}}^{(3)+}_{-}&{\widehat{\mathbf{A}}}^{(3)+}_{(2)+}&{\widehat{\mathbf{A}}}^{(3)+}_{(2)-}&{\widehat{\mathbf{A}}}^{(3)+}_{(3)+}&{\widehat{\mathbf{A}}}^{(3)+}_{(3)-}&\ddots\\ {\widehat{\mathbf{A}}}^{(3)-}_{(0)}&{\widehat{\mathbf{A}}}^{(3)-}_{+}&{\widehat{\mathbf{A}}}^{(3)-}_{-}&{\widehat{\mathbf{A}}}^{(3)-}_{(2)+}&{\widehat{\mathbf{A}}}^{(3)-}_{(2)-}&{\widehat{\mathbf{A}}}^{(3)-}_{(3)+}&{\widehat{\mathbf{A}}}^{(3)-}_{(3)-}&\ddots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\ddots&\ddots\end{pmatrix}\begin{pmatrix}1\\ \mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\\ \mathbf{x}^{(3)+}\\ \mathbf{x}^{(3)-}\\ \vdots\end{pmatrix}$
@@
(243) $\left|\langle\mathbf{x}^{(m)}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle\right|$ $\leq\mathcal{O}(a,b),\quad m\geq 4$
(244) $\left|\langle\mathbf{x}^{(m)}(\tau){\mathbf{x}^{(2)}(0)}^{\mathsf{T}}\rangle\right|$ $\leq\mathcal{O}(a,b),\quad m\geq 3,$
@@
(246) $\left|{{}\widehat{\mathbf{A}}^{n}}^{(m)}\right|$ $\leq\mathcal{O}(a,b),\quad m\geq 4$
(247) $\left|{{}\widehat{\mathbf{A}}^{n}}^{(m)}_{(2)}\right|$ $\leq\mathcal{O}(a,b),\quad m\geq 3.$
@@
Again, the base case $n=1$ for Eq. (249) follows from the aforementioned conditions. To carry out the proof by induction for Eq. (248)–(249) assuming the base case $n=1$, we need to use:
@@
(251) ${{}\widehat{\mathbf{A}}^{n}}^{(i)}_{(i)}={{{}\widehat{\mathbf{A}}}^{(i)}_{(i)}}^{n}+\mathcal{O}((a,b)^{2}).$

### 11 Jan 2024 v6

@@ 6.2. One-dimensional case
Apparently, there are 6 parameters ($A$, $D$, $a$, $b$, $\mu$, $\nu$); however, there are actually only 5 (quadratic drift has only 3 coefficients). Thus, if $\mu$ and $\nu>0$ are the actual mean and variance of $x$, then given these together with $D>0$ and $b$, the possible values of the pair $(A,a)$ must lie along a one-dimensional manifold. Thus in Eq. (81), $\mathbf{C}$ should be interpreted as a “nominal” value not necessarily corresponding to the actual covariance matrix. In particular, we will take it to be the value obeying the Lyapunov equation for the linear part of the dynamics (Eq. (3)).
@@
$\qquad{}+a^{2}\mathcal{O}(1){He}_{3}(x)+a^{3}\mathcal{O}(1){He}_{4}(x)+\cdots,$
@@
$\qquad{}+(2a+a^{2}\mathcal{O}(a,b)){He}_{3}(x)+a^{2}\mathcal{O}(1){He}_{4}(x)+\cdots,$
@@
We now present some numerical simulations to confirm the above results. We simulate a one-dimensional stochastic process where $\langle\dot{x}\mid x\rangle$ follows Eq. (123) with $x\to x+a$ and inhomogeneous diffusion $(1+bx/2)^{2}$. We use an Euler–Maruyama discretization with time-step $\Delta t=0.05$ and trajectory length $10^{6}$ time-units. Although corrections to the third-order covariance functions are only cubic order in $a,b$, they turn out to be quite large, as shown in Fig. 5 (the exception being when $a=0$). The second-order covariance function is also affected, although less so (Fig. 6), but the deviation is still larger than what one might expect for a quadratic correction. However, the deviations are limited to approximate multiplication by a constant, and thus nonlinearity cannot be effectively discerned from the second-order covariance functions, as expected. For larger values of $|a|$ ($\sim$ 0.3), the covariance functions are nowhere near the theoretically predicted values (not shown).
@@ 6.5. Integrated variables
We change coordinates to ${x^{\prime}}^{i}:=f_{1}^{i}(\mathbf{x})$, where
@@
6.6. Second-order covariance functions to quadratic order
We may calculate second-order covariance functions to order $a\mathcal{O}(a,b)$, with $a^{2}\mathcal{O}((a,b)^{2})$ corrections. We may consider what happens if second-order angular momenta are $a^{2}\mathcal{O}((a,b)^{2})$ and third-order angular momenta are $a\mathcal{O}((a,b)^{2})$. We again change coordinates to ${x^{\prime}}^{i}:=f_{1}^{i}(\mathbf{x})$. We have:
(169) $L({x^{\prime}}^{i},{x^{\prime}}^{j})$ $=a^{2}\mathcal{O}((a,b)^{2}),$
(170) $L({x^{\prime}}^{i}{x^{\prime}}^{j},{x^{\prime}}^{k})$ $=a\mathcal{O}((a,b)^{2}),$
and therefore:
(171) $\langle{x^{\prime}}^{i}(0){x^{\prime}}^{j}(\tau)\rangle-\langle{x^{\prime}}^{i}(\tau){x^{\prime}}^{j}(0)\rangle$ $=a^{2}\mathcal{O}((a,b)^{2}),$
(172) $\langle{x^{\prime}}^{i}(0){x^{\prime}}^{j}(0){x^{\prime}}^{k}(\tau)\rangle-\langle{x^{\prime}}^{i}(\tau){x^{\prime}}^{j}(\tau){x^{\prime}}^{k}(0)\rangle$ $=a\mathcal{O}((a,b)^{2}).$
Switching back to the original coordinates:
(173) $x^{i}=\mathcal{N}^{i}_{j}\left[{{x^{\prime}}^{j}-\alpha^{j}_{kl}\mathcal{O}(1)\left({{x^{\prime}}^{k}{x^{\prime}}^{l}-{C^{\prime}}^{kl}}\right)+a^{2}\mathcal{O}(1){He}_{3}^{\prime}(\mathbf{x}^{\prime})+a^{3}\mathcal{O}(1){He}_{4}^{\prime}(\mathbf{x}^{\prime})+\cdots}\right],$
where $\mathcal{N}^{i}_{j}=\delta^{i}_{j}+a^{2}\mathcal{O}(1)$ is a normalization factor, $\mathbf{C}^{\prime}:=\langle\mathbf{x}^{\prime}{\mathbf{x}^{\prime}}^{\mathsf{T}}\rangle$, and ${He}_{n}^{\prime}$ denotes the multidimensional Hermite polynomial defined with respect to covariance matrix $\mathbf{C}^{\prime}$. We then have:
(174) $\langle x^{i}(0)x^{j}(\tau)\rangle-\langle x^{i}(\tau)x^{j}(0)\rangle=a^{2}\mathcal{O}((a,b)^{2}),$
i.e., the second-order covariance functions are symmetric to quadratic order. A similar result holds for integrated variables.
@@ 7.3. Inhomogeneous diffusion
(233) $\left\langle{\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}\mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\end{pmatrix}\mid\mathbf{x}}\right\rangle=\begin{pmatrix}\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&&\\ \mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&&\\ \widehat{\mathbf{A}}^{(2)+}_{+}&\widehat{\mathbf{A}}^{(2)+}_{-}&\widehat{\mathbf{A}}^{(2)+}_{(2)+}&\widehat{\mathbf{A}}^{(2)+}_{(2)-}\\ \widehat{\mathbf{A}}^{(2)-}_{+}&\widehat{\mathbf{A}}^{(2)-}_{-}&\widehat{\mathbf{A}}^{(2)-}_{(2)+}&\widehat{\mathbf{A}}^{(2)-}_{(2)-}\end{pmatrix}\begin{pmatrix}\mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\end{pmatrix}.$
We again assume, as throughout this section, that $\mathbf{D}^{+-}=\mathbf{0}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle=\mathbf{0}$. To have time-reversibility of the covariance functions, we first need that:
@@
(237) ${{}\widehat{\mathbf{A}}^{n}}^{(2)\pm}_{\mp}\langle\mathbf{x}^{\mp}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{(2)\pm}_{(2)\mp}\langle\mathbf{x}^{(2)\mp}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle$ $=-\langle\mathbf{x}^{(2)\pm}{\mathbf{x}^{\pm}}^{\mathsf{T}}\rangle{{\mathbf{A}^{n}}^{\mp}_{\pm}}^{\mathsf{T}}.$
@@
(239) $\left\langle{\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\\ \mathbf{x}^{+}\Delta y\\ \mathbf{x}^{-}\Delta y\end{pmatrix}\mid\mathbf{x}}\right\rangle=\begin{pmatrix}0&&&&&&\\ &\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&&&&\\ &\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&&&&\\ &\widehat{\mathbf{A}}^{(2)+}_{+}&\widehat{\mathbf{A}}^{(2)+}_{-}&\widehat{\mathbf{A}}^{(2)+}_{(2)+}&\widehat{\mathbf{A}}^{(2)+}_{(2)-}&&\\ &\widehat{\mathbf{A}}^{(2)-}_{+}&\widehat{\mathbf{A}}^{(2)-}_{-}&\widehat{\mathbf{A}}^{(2)-}_{(2)+}&\widehat{\mathbf{A}}^{(2)-}_{(2)-}&&\\ \widehat{\mathbf{A}}^{y(1)+}_{(0)}&\widehat{\mathbf{A}}^{y(1)+}_{+}&\widehat{\mathbf{A}}^{y(1)+}_{-}&\widehat{\mathbf{A}}^{y(1)+}_{(2)+}&\widehat{\mathbf{A}}^{y(1)+}_{(2)-}&\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}\\ \widehat{\mathbf{A}}^{y(1)-}_{(0)}&\widehat{\mathbf{A}}^{y(1)-}_{+}&\widehat{\mathbf{A}}^{y(1)-}_{-}&\widehat{\mathbf{A}}^{y(1)-}_{(2)+}&\widehat{\mathbf{A}}^{y(1)-}_{(2)-}&\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}\end{pmatrix}\begin{pmatrix}1\\ \mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\\ \mathbf{x}^{+}\Delta y\\ \mathbf{x}^{-}\Delta y\end{pmatrix}.$
@@
(242) ${{}\widehat{\mathbf{A}}^{n}}^{y(1)+}_{-}\langle\mathbf{x}^{-}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle+{{}\widehat{\mathbf{A}}^{n}}^{y(1)+}_{(2)-}\langle\mathbf{x}^{(2)-}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle$ $=\langle\mathbf{x}^{+}{\mathbf{x}^{+}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)-}_{+}}^{\mathsf{T}}+\langle\mathbf{x}^{+}{\mathbf{x}^{(2)+}}^{\mathsf{T}}\rangle{{{}\widehat{\mathbf{A}}^{n}}^{y(1)-}_{(2)+}}^{\mathsf{T}}.$
Again, the base case $n=1$ of Eq. (242) is a consequence of Eqs. (234) and (240), while Eq. (241)–(242) can be shown by mathematical induction on $n$.
@@
(243) $\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\\ \langle\mathbf{x}^{(2)+}\Delta y\rangle\\ \langle\mathbf{x}^{(2)-}\Delta y\rangle\end{pmatrix}=\begin{pmatrix}0&&&&\\ {\widehat{\mathbf{A}}}^{y(1)+}_{(0)}&\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&&\\ {\widehat{\mathbf{A}}}^{y(1)-}_{(0)}&\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&&\\ {\widehat{\mathbf{A}}}^{y(2)+}_{(0)}&\widehat{\mathbf{A}}^{(2)+}_{+}&\widehat{\mathbf{A}}^{(2)+}_{-}&\widehat{\mathbf{A}}^{(2)+}_{(2)+}&\widehat{\mathbf{A}}^{(2)+}_{(2)-}\\ {\widehat{\mathbf{A}}}^{y(2)-}_{(0)}&\widehat{\mathbf{A}}^{(2)-}_{+}&\widehat{\mathbf{A}}^{(2)-}_{-}&\widehat{\mathbf{A}}^{(2)-}_{(2)+}&\widehat{\mathbf{A}}^{(2)-}_{(2)-}\end{pmatrix}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\\ \langle\mathbf{x}^{(2)+}\Delta y\rangle\\ \langle\mathbf{x}^{(2)-}\Delta y\rangle\end{pmatrix}.$
@@
Once again, the case $n=1$ for Eq. (245) is a consequence of Eqs. (234) and (240), while Eq. (244)–(245) can be shown assuming the case $n=1$.
@@ 7.4. Nonlinear drift
(246) $\left\langle{\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\\ \mathbf{x}^{(3)+}\\ \mathbf{x}^{(3)-}\\ \vdots\end{pmatrix}\mid\mathbf{x}}\right\rangle=\begin{pmatrix}0&&&&&&&\\ &\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&{\widehat{\mathbf{A}}}^{+}_{(2)+}&{\widehat{\mathbf{A}}}^{+}_{(2)-}&&&\\ &\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&{\widehat{\mathbf{A}}}^{-}_{(2)+}&{\widehat{\mathbf{A}}}^{-}_{(2)-}&&&\\ &{\widehat{\mathbf{A}}}^{(2)+}_{+}&{\widehat{\mathbf{A}}}^{(2)+}_{-}&{\widehat{\mathbf{A}}}^{(2)+}_{(2)+}&{\widehat{\mathbf{A}}}^{(2)+}_{(2)-}&{\widehat{\mathbf{A}}}^{(2)+}_{(3)+}&{\widehat{\mathbf{A}}}^{(2)+}_{(3)-}&\\ &{\widehat{\mathbf{A}}}^{(2)-}_{+}&{\widehat{\mathbf{A}}}^{(2)-}_{-}&{\widehat{\mathbf{A}}}^{(2)-}_{(2)+}&{\widehat{\mathbf{A}}}^{(2)-}_{(2)-}&{\widehat{\mathbf{A}}}^{(2)-}_{(3)+}&{\widehat{\mathbf{A}}}^{(2)-}_{(3)-}&\\ {\widehat{\mathbf{A}}}^{(3)+}_{(0)}&{\widehat{\mathbf{A}}}^{(3)+}_{+}&{\widehat{\mathbf{A}}}^{(3)+}_{-}&{\widehat{\mathbf{A}}}^{(3)+}_{(2)+}&{\widehat{\mathbf{A}}}^{(3)+}_{(2)-}&{\widehat{\mathbf{A}}}^{(3)+}_{(3)+}&{\widehat{\mathbf{A}}}^{(3)+}_{(3)-}&\ddots\\ {\widehat{\mathbf{A}}}^{(3)-}_{(0)}&{\widehat{\mathbf{A}}}^{(3)-}_{+}&{\widehat{\mathbf{A}}}^{(3)-}_{-}&{\widehat{\mathbf{A}}}^{(3)-}_{(2)+}&{\widehat{\mathbf{A}}}^{(3)-}_{(2)-}&{\widehat{\mathbf{A}}}^{(3)-}_{(3)+}&{\widehat{\mathbf{A}}}^{(3)-}_{(3)-}&\ddots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\ddots&\ddots\end{pmatrix}\begin{pmatrix}1\\ \mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \mathbf{x}^{(2)+}\\ \mathbf{x}^{(2)-}\\ \mathbf{x}^{(3)+}\\ \mathbf{x}^{(3)-}\\ \vdots\end{pmatrix}.$
@@
(249) $\langle\mathbf{x}^{(4)}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ $=\mathcal{O}(a,b),$
@@
(251) ${{}\widehat{\mathbf{A}}^{n}}^{(4)}$ $=\mathcal{O}(a,b).$
@@
Again, the base case $n=1$ for Eq. (253) follows from Eqs. (234)–(235). To carry out the proof by induction for Eq. (252)–(253) assuming the base case $n=1$, we need to use, for $i\geq 1$:
@@
(255) ${{}\widehat{\mathbf{A}}^{n}}^{(i)}_{(i)}={{{}\widehat{\mathbf{A}}}^{(i)}_{(i)}}^{n}+a\mathcal{O}(a,b).$